OpenStudy (anonymous):

What is the value of tan1o tan2o tan3o…………. tan 87o tan88o tan89o?

OpenStudy (anonymous):

tan(1+2+3+........89)

OpenStudy (anonymous):

@Auxuris

ganeshie8 (ganeshie8):

hint : \(\tan(\theta) = \cot(90 - \theta) = \dfrac{1}{\tan(90-\theta )}\)

ganeshie8 (ganeshie8):

\[\large \tan 1 \tan 2 \tan 3 \cdots \tan 87 \tan 88 \tan 89\]

ganeshie8 (ganeshie8):

focus on first and last terms \[\large \color{Red}{\tan 1} \tan 2 \tan 3 \cdots \tan 87 \tan 88 \color{red}{\tan 89}\]

OpenStudy (anonymous):

tan1=cot(90-1)=cot89

ganeshie8 (ganeshie8):

yes which equals ?

OpenStudy (anonymous):

cot89.tan89=1

ganeshie8 (ganeshie8):

tan1=cot(90-1)=cot89 = 1/tan89 right ?

ganeshie8 (ganeshie8):

yep !

ganeshie8 (ganeshie8):

so basically we can pair up and cancel everything except the `middle term` any idea what the middle term would be ?

OpenStudy (anonymous):

hey can we do yesterdays problem

OpenStudy (anonymous):

tan45=1

ganeshie8 (ganeshie8):

il msg u once i figure out how to solve yesterday's problem :)

OpenStudy (anonymous):

using cos and sin terms instead of tan2x-sec2x=1